OpenStudy (anonymous):

If the midpoint between (x, 3) and (9, 9) is (7, 6), find the value of x. A. -2 B. 23 C. 5 D. 9

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(x\quad ,&3)\quad &(9\quad ,&9) \end{array}\\ \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\implies \left(\cfrac{9 + x}{2}\quad ,\quad \cfrac{9 + 3}{2} \right)=(7,6)\qquad thus\\ \quad \\ \cfrac{9 + x}{2}=7\)

OpenStudy (jdoe0001):

solve for "x" to find "x"